OpenStudy (anonymous):

@satellite73 @ganeshie8

OpenStudy (shadowlegendx):

What is your question?

OpenStudy (anonymous):

Solve. \[\log_{4} \sqrt{8} =\]

OpenStudy (anonymous):

there is supposed to be an X after the equals sighn

OpenStudy (anonymous):

x is not important this is asking 'what power do you raise 4 to, to give you 8 i.e. is asking "solve for \(x\), \(4^x=8\)

OpenStudy (anonymous):

oh square root of 8 sorry \[4^x=\sqrt{8}\]

OpenStudy (anonymous):

I got 2.82842712 for the square root of 8

OpenStudy (anonymous):

they want you to recognize both as powers of 2 \[4^x=2^{2x}\] \[\sqrt{8}=\sqrt{2^3}=2^{\frac{3}{2}}\]

OpenStudy (anonymous):

this is not a calculator exercise

OpenStudy (anonymous):

\[\large 2^{2x}=2^{\frac{3}{2}}\iff 2x=\frac{3}{2}\]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

lol no you have to solve \[\large 2x=\frac{3}{2}\] for \(x\)

OpenStudy (anonymous):

ok how?

OpenStudy (shadowlegendx):

I would multiply both sides by 2 to make them wholes

OpenStudy (shadowlegendx):

4x = 3

OpenStudy (anonymous):

if you wanted to use a calculator to solve this, you can you type in \[\log(\sqrt8)\div\log(4)\] and you will get it

OpenStudy (shadowlegendx):

Oh, haha c;

OpenStudy (anonymous):

@ShadowLegendX no ! divide by 2 and get \[x=\frac{3}{4}\]

OpenStudy (anonymous):

thank you soo much I appreciate it

OpenStudy (anonymous):

yw

OpenStudy (shadowlegendx):

Ok, good job @satellite73